UARS Weekly Status Report

27 April 2001




********************************************************************

                UARS WEEKLY STATUS REPORT - 27 APR 2001

********************************************************************

***********
SOLSTICE
***********

Nothing received.

***********
SUSIM
***********

SUSIM was switched off for 18-22 April because of spacecraft power
constraints.  Otherwise SUSIM performed its daily solar scans this
past week during periods of predicted UARS telemetry.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

     At the present time, PEM energy deposition from its spectrometers is not 
being generated.  Recent data configuration changes have affected PEM software 
causing multiple failures.  These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update 
processing code with a patch delivery.  The delivery patch may include changes 
which would allow the PEM software to execute on a DEC Alpha machine.  GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).

     PEM has examined its processed files of the rehost machine and has found 
that no differences exist in the Level 1 and Level 2 output files which were 
generated the rehost when the same input data was used on CDHF and CDHFB.
PEM is now confident that output generated by PEM 3.32, Jobs 1 & 2, on the 
rehost system is valid and correct.

     The PEM team is working on replacing the PEM:: machine with a DEC Alpha 
machine to be called PEM2K::.  Tests of the NRT data transfer from CDHF2:: 
through PEM2K:: to the unix computer running the PEM GSE software show no 
difference between equivalent NRT files transferred from PEM::.  Thus, NRT 
data from CDHF2:: is captured correctly on PEM2K::.  Tests of the NRT data 
transfer from CDHFB:: were not successful from both PEM:: and PEM2K::.  
Results of the test of NRT data transfers from the rehost computer were
successful for individual files which exist on CDHFB::.  PEM will participate
in the NRT data flow to capture NRT data from CDHFB:: when specified by GSFC.

     PEM has been working with GSFC on allowing commanding of the PEM
instrument by creating DAPs on the PEM2K machine and submitting them to the 
CMS.  So far, this has not been achieved.  It was discovered that PEM2K is 
running Openwindows version 7.2-1 which operates differently across the 
network than Openwindows version 7.1-2 (the CDHFB operating system).  Compaq 
is investigating to determine a solution.  In the mean time, Steve Bove made
some suggestions which seem to have improved the PEM2K:: to UCMS:: image
execution ability; however, the CMS has had troubles recently and it appears
that the UCMS:: definitions for PEM2K:: must not be correct on the UCMS:: node.
PEM is waiting on a reply back from GSFC.

    Updated HEPS calibration parameters have been finalized and are ready for 
delivery.  Updated MEPS quality calibration files have been finalized and are 
ready for delivery.  Level 3 software errors, discovered by Dr. Dean Pesnell, 
have been fixed; his suggestions have been incorporated with in the PEM 
production code.  Both updated HEPS and MEPS calibration data, along with the 
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 30 May 2001.

***********
CDHF
***********

HIGHLIGHTS
----------
- All level 0 data from the power outage time frame are available.

UARS CDHF REHOST
----------------
- Additional 2 weeks have been requested to confirm "puts" to silo. Once this is
done we are able to go operational.
We are anxious to go operational on the REHOST due to resource requirements
needed to run two clusters.       
                  
***********
FOT
***********

                         REPORTING PERIOD

This report covers 14 April 2001 (Orbit 52321, GMT Day 104) through 20 April 
2001 (Orbit 52425, GMT Day 110).

                       SPACECRAFT OPERATION

The observatory is now in Forward Flight (Southern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


                  INSTRUMENT OPERATIONAL CHANGES


 Instrument     Time       Orbit     Comment          
 HALOE        106/0241     52457     Power OFF
 SUSIM        109/0006     52500     Power OFF
 ACRIM        109/0006     52500     Power OFF
 PEM          109/0106     52501     Power OFF
 HRDI         109/2313     52515     Low Power          

A Reverse-to-Forward Yaw Around maneuver (number 98) was successfully 
executed in Orbit 52510 beginning 109/16:01:03 and completing at 109/16:38:00 
on 19 April 2001.  The SOLSTICE instrument was in day only and all other 
instruments were in their off or standby mode (including HRDI and PEM).  Due 
to the SA position at 269 degrees, the effective night length of the 
spacecraft after the maneuver was approximately 55.9 minutes.  Post maneuver 
analysis of data indicate that during the Yaw Around, the end-of-night load 
bus voltage was 23.90 volts and the average depths of discharge (DODs) on 
Batteries 2 and 3 were approximately 22.8 and 23.0 percent respectively.  The 
next (forward-to-reverse) YAW is scheduled for 30 May 2001 (DOY 150).

The beta angle ranged between the angles of -24.2 degrees to +1.4 degrees for 
this report period.  The beta angle is now increasing towards a maximum angle 
of 74 degrees on 09 May 2001 (DOY 129).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and 
Battery 3 of 1.3 and 1.6 degrees C.  Additional battery performance data for 
each battery on 13 April 2001 and 20 April 2001 are provided in Appendix D.

The clock error ranged from –13.2 to 13.4 msec during this report period.  
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion 
of ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits 
occurrences are listed in Appendix B.

•     OUT-OF-LIMITS  =  02
      ALERTS         =  00

The following real-time command sheets were executed for the listed 
operational element(s) during this report period (see Appendix A).

•     FOT            =  18
•     HALOE          =  01
•     SUSIM          =  01
•     ACRIM          =  01
•     PEM            =  01
•     HRDI           =  01


UARS Anomaly Reports initiated during this report period are listed in 
Appendix E and summarized below.

•     FOT            =  02

One (1) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
∑ Transponder A:  00/316  - 2287.496493
∑ Transponder B:  01/087  - 2287.499152

                      GROUND SYSTEM OPERATION

Two (2) operational support problems occurred this period.  Anomaly Reports 
are included in Appendix E.

The data loss calculations supplied by Data Capture Facility (DCF) from 
mission start through UTC Day 01/067 show the observatory data loss to be 300 
hours, 3 minutes, 5.464 seconds. This shows an increase of 1 minute and 
52.704 seconds since the last report period. This is a 0.39 percent data 
loss, which equals a 99.61 percent data capture for the mission.

∑ 15.424 seconds on Day 107 due to a Data Hit,
∑ 1 minute and 37.280 seconds on Day 108 due to a Data Hit.

                     OTHER SIGNIFICANT EVENTS

                     UARS SATELLITE OPERATIONS

UARS is using both the Fine Sun Sensor (FSS) and the Three Axis Magnetometer 
(TAM) as inputs to UFLTR this week.  The FOT is pleased with their 
performance, but continues to make adjustments to the TAM alignment to 
improve its performance and UARS’ attitude.

The FSW patch that will improve the magnetic field model from a 5th order to 
7th order solution has been completed.  The patch requires one further test 
next week before the planned uplink date of 25 April 2001.

                           MEETINGS

The FOT attended a UARS TAC RS Replacement Meeting on Wednesday, April 18, 
2001.The FOT attended the UARS SERB Meeting on Wednesday, April 18, 2001.


                   FUTURE SIGNIFICANT EVENTS

The next UARS Yaw slew is scheduled for 30 May 2001 (DOY 150).  This maneuver 
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).


Attachments:  Appendix A,  Real-time Commands Sheets
              Appendix B,  Out-Of-Limits & ALERT Occurrences
              Appendix C,  Clock & Solar Array Rate Corrections
              Appendix D,  UARS Battery Performance Data
              Appendix E,  UARS Anomaly Reports

                         APPENDIX A
                  REAL-TIME COMMAND SHEETS
                   Week ending 20 April 2001

ORBIT    AOS        NO.      SUBSYS     REASON
-----    ---        ---      ------     ------
52430    104/0750   01-231   NBTR       NOOPS D-104
52437    104/1805   01-232   NBTR       NOOPS D-105
52443    105/0420   01-233   OBC/CLK    Clk Adj 0.00 to -20.6 msec/day
52444    105/0646   01-232   NBTR       NOOPS D-105
52452    105/1805   01-234   NBTR       NOOPS D-106
52457    106/0241   01-235   HALOE      Power OFF
52459    106/0605   01-236   OBC/CLK    Clk Adj -20.6 to 0.00 msec/day
52460    106/0754   01-234   NBTR       NOOPS D-106
52465    106/1443   01-237   OBC/CLK    Clk Adj 0.00 to 20.6 msec/day
52478    107/1207   01-238   OBC/CLK    Clk Adj 20.6 to 0.00 msec/day
52483    107/2010   01-239   OBC/TAM    MAGMOD Time Update
52488    108/0424   01-240   OBC        Clk Adj fr 0.00 to -20.6 msec/day
52493    108/1211   01-241   NBTR       PBK
52500    109/0006   01-242   SUSIM      Power OFF
52500    109/0006   01-243   ACRIM      Power OFF
52501    109/0106   01-244   PEM        Power OFF
52501    109/0106   01-245   SOLSTICE   TO DAY OPS ONLY
52508    109/1209   01-246   OBC/CLK    Clk Adj -20.6 to 0.00 msec/day
52510    109/1529   01-250   OBC/ACS    YAW MANEUVER
52510    109/1529   01-249   OBC/CLK    Clk Adj 0.00 to 20.6 msec/day
52510    109/1529   01-247   OBC/TAM    TAM TO MONITOR MODE
52512    109/1842   01-248   OBC/TAM    TAM TO UPDATE MODE
52515    109/2313   01-251   HRDI       Low Power
52525    110/1617   01-252   OBC        Clk Adj fr 20.6 to 0.00 msec/day 
________________________________________________________________________

                          APPENDIX B
            OUT-OF-LIMITS & ALERTS OCCURRENCES
                 Week ending 20 April 2001

The following Out-Of-Limits occurrences were detected:

MNEMONIC        STATE     DESCRIPTION                  ORBITS
--------        -----     -----------                  ------
HRIFESURVTMP    Y-HI      HRDI IFE Survival Htr Temp   52517
HRRADTEMP       Y-HI      HRDI Radiator Temp           52517


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION     ORBITS
--------     -----------     ------
-None this report period

_____________________________________________________________________________

                         APPENDIX C
          CLOCK & SOLAR ARRAY RATE CORRECTIONS
                 Week ending 20 April 2001

               CLOCK DRIFT RATE ADJUSTMENTS

ORBIT     TIME         CLOCK RATE CHANGE
-----     -----        -----------------

52443     105/0420     Clk Adj fr 0.00 to -20.6 msec/day
52459     106/0605     Clk Adj fr -20.6 to 0.00 msec/day
52465     106/1443     Clk Adj fr 0.00 to 20.6 msec/day
52478     107/1207     Clk Adj fr 20.6 to 0.00 msec/day
52488     108/0424     Clk Adj fr 0.00 to -20.6 msec/day
52508     109/1209     Clk Adj fr -20.6 to 0.00 msec/day
52510     109/1529     Clk Adj fr 0.00 to 20.6 msec/day
52525     110/1617     Clk Adj fr 20.6 to 0.00 msec/day     
________________________________________________________________________

                       APPENDIX D
              UARS BATTERY PERFORMANCE DATA
                Week ending 20 April 2001


     13 Apr 2001 - GMT Day 103 Beta = 27.8 deg, SA Pos = 269 deg

     DIFF V (mV)   TEMP (C)     CURR (amp)    EON V   AVE C/D   AVE
BAT  MAX/MIN       MAX/MIN      MAX/MIN       MIN     RATIO     DOD(%)
---  -----------   --------     ----------    ------  -------   -----
1    +67.2/+44.8   -3.05/-4.31  +0.0/+0.0     13.0    0.000      0.0
2    +11.2/-100.8  +6.44/+5.15  +29.6/-15.6   24.0    1.040      21.8
3    +22.4/-67.2   +5.47/+3.87  +28.8/-16.0   23.8    1.030      21.7


     20 Apr 2001 - GMT Day 110 Beta = 01.4 deg, SA Pos = 269 deg

     DIFF V (mV)   TEMP (C)     CURR (amp)     EON V    AVE C/D  AVE
BAT  MAX/MIN       MAX/MIN      MAX/MIN        MIN      RATIO    DOD(%)
---  -----------   --------     ----------     ------   -------  -----
1   +621.6/+280.0  -3.68/-4.94  +0.0/+0.0      12.3     0.000    0.0
2   +22.4/-67.2    +5.15/+3.87  +28.8/-12.8    25.1     1.020    20.5
3   +22.4/-67.2    +3.87/+2.60  +28.0/-12.8    25.0     1.020    20.0


     V/T LEVEL CHANGES
ORB#            TIME          LVL FR-TO
----            ----          ---------
-no changes     V/T 5
_____________________________________________________________________________
             
                  APPENDIX E
                 UARS ANOMALY REPORTS
               Week ending 20 April 2001

              NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
01-019     FOT        52501     109/0146   14170   Generic Late Acquisition
01-019     FOT        52506     109/0944   14170   Generic Late Acquisition

            ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
-None this report period

           ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
-None this report period

                    OPEN ANOMALY REPORTS
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
01-019     FOT        see list  Apr 01     14170   Generic Late Acquisition


	 	SEVNTFW01/16